OpenStudy (jiteshmeghwal9):

If sum of the roots is 'p' & the sum of their squares is \[q^2/] , then the equation is ;-

OpenStudy (jiteshmeghwal9):

\[Ans. x^2-px+(p^2-q^2)/2=0\]

OpenStudy (matheducatormcg):

weird. where's the question from? (what class I mean)

OpenStudy (foolaroundmath):

Let the roots be \(\alpha \) and \(\beta\). Given : \(\alpha+ \beta = p \) \(\alpha ^{2} + \beta ^{2} = q^{2} \) \( (\alpha + \beta)^{2} = \alpha ^{2} + \beta ^{2} + 2 \alpha \beta\) \(\Rightarrow p^{2} = q^{2} + 2 \alpha \beta\) \(\Rightarrow \alpha \beta = (p^{2} - q^{2})/2\) So, sum of roots = p, product of roots = \((p^{2}-q^{2})/2\). Hence the quadratic equation is \(x^{2} - px + (p^{2} - q^{2})/2 = 0 \)

OpenStudy (matheducatormcg):

word

OpenStudy (jiteshmeghwal9):

Thanx! @FoolAroundMath